Circular No. 7267

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 S4 (LINEAR) Another unusual-moving object discovered by LINEAR and reported

as asteroidal, subsequently posted on the NEO Confirmation Page, has been noted to be cometary in appearance by D. Durig (Sewanee, TN, 0.3-m f/7 reflector + CCD; coma diameter ∼ 10′′; tail ∼ 20′′–25′′ long in p.a. 200^{o}– 220^{o}) and by J. Tichá and M. Tichý (Kleť, 0.57-m f/5.2 reflector + CCD; comet diffuse with 8′′ coma and tail > 10′′ long in p.a. 245^{o}). The discovery observation is given below, and additional astrometric observations appear on MPEC 1999-T02:

1999 UT  α_{2000}  δ_{2000}  m_{2}
Sept.27.39772  5^{h}03^{m}47.^{s}42  +34^{o}16′45′′.3  17.3

The following preliminary parabolic orbital elements (from MPEC 1999-T02) suggest that this comet might become a naked-eye object next July:

T = 2000 July 18.288 TT  ω = 152.^{o}651
Ω = 83.387 }2000.0
q = 0.71953 AU  i = 149.934
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 29 5^{h}02.^{m}95 +34^{o}24.^{′}1 3.851 4.254 107.^{o}0 13.^{o}0 16.2
Oct. 4 4 59.84 +34 47.4 3.716 4.200 112.4 12.7 16.1
9 4 55.95 +35 10.8 3.584 4.146 117.9 12.3 15.9
14 4 51.21 +35 33.9 3.455 4.092 123.5 11.7 15.8
19 4 45.56 +35 56.1 3.332 4.037 129.2 11.0 15.7

GM SAGITTARII P. A. Charles and T. Shahbaz, University of Oxford; and T. Geballe,

Gemini Observatory, report: “We have obtained a K-band (1.9–2.5-μm) low-resolution (2.5-nm) spectrum of the x-ray transient GM Sgr = XTE J1819–254 (IAUC 7253) on Sept. 17.22 UT with the U.K. Infrared Telescope (+ CGS4). We estimate K = 12.5 ± 0.3. Our spectrum shows a blue continuum and broad emission lines of He I (2.06 μm; EW = 2.1 nm) and Brγ (EW = 1.4 nm), characteristic of low-mass x-ray binaries. Both profiles have a FWZI of ∼ 5900 km/s. There is clear evidence for an extended blue wing in the Brγ profile, which may be related to a highvelocity wind component. Further spectroscopic monitoring during this outburst is encouraged.”

1999 October 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green